Find the number of 15-centimeter square tiles required to cover a rectangular floor 5.4 m long and 4.05 m wide.
Answer by mananth(16946)

Number of tiles required = area of floor/area of tile
area of floor = 5.4*4.05=21.87 m^2
Area of tile = 0.15 *0.15 = 0.0625 m^2
21.87/0.0625=349.92
= 350 tiles